In 3 bowling games, Alice scores 139, 143 and 144. What score will Alice need in a fourth game in order to have an average score of 145 for all four games?
Answer by Horlaye(27)

Let p represent the fourth score
(add all scores/occurrence = average)
(139+143+144+p)/4 = 145
(426+p)/4 = 145/1
Cross multiply
. 426+p = 4(145)
426+p = 580
Collect like term
p = 580-426
p = 154
Since p represent the fourth score
Therefore Alice need to score 154 in her fourth game to get the average score of 145